How many molecules of C9H20 would you have if you have 1.91 mol?
Answer #1

1 mole of any substance contains 6.023*10^23 molecules

1 mole of C9H20 contains 6.023*10^23 molecules

1.91 mole of C9H20 contains = 6.023*10^23molecules *1.91mole/1mole = 1.15*10^24 molecules

1.15*10^24 molecules of C6H20 >>>>>answer